# Moin et al

### v2

@@ 4 Research design and methodology
We conducted research in two steps: 1) We interviewed four selected domain experts to gain some insight into the problem domain and validate the interview questions. To this aim, we deployed the Qualitative Surveys (Interview Studies) empirical research method 21. 2) We adopted the quantitative empirical research method, called Questionnaire Surveys 22 and carried out an online survey with 61 participants.
@@
We interviewed four experts from the authors’ networks from May to June 2021. The interviewee selection was based on convenience sampling (i.e., not random). The interviewee profiles were as described in Table 2. The numbers in parentheses denote the frequencies. The interviews were semi-structured with open questions.
Table 2: Summary of the demographic information of the interview partners
@@
We carried out a survey study from July to September 2021 to answer RQ1 and RQ2 above. The survey questionnaire was offered through a link (URL) as an online [^5] We used the open-source LimeSurvey software 23 on our own server., self-administered survey, with the option of fully anonymous participation in the study. However, we collected the IP addresses to prevent any possible redundant participation. The questionnaire had four sections and a total of 25 questions. The order of the questions in each section and the order of the choices in the case of multiple-choice questions would be set on a random basis for each participant. We had a total of 121 participants, out of which 60 participants did not answer the questionnaire at all. Therefore, we took the results of the remaining 61 participants into account. The selection process for the invitation of the subject matter experts to participate in this study was again based on convenience sampling (i.e., not random) through the authors’ networks, for example, by direct invitation of peers via email and sharing the URL on LinkedIn. Part of the participants’ demographic information is summarized in Table 3 (the numbers in parentheses denote the frequencies). The average participation time was 14 minutes, whereas the median was 11 minutes. All questions were optional to answer except for the two questions regarding the consent of the survey participants with respect to their anonymity and receiving a pre-print of the study in the future.
Table 3: Summary of the survey participants’ demographic information
@@
Table 4: Summary of the demographic information of the interview partners in the second round of expert interviews
@@
5 Enhancing architecture frameworks
Existing software system architecture frameworks fall short of supporting a holistic architecture description for ML-enabled systems. Some members of the Software Engineering (SE) community may argue that the ML-related functionality might be specified using existing functional, behavioral, or logical model kinds, and perhaps the structure of ML models might be specified using existing structural model kinds. However, looking at the reality and state of practice in ML, data science, and data engineering, as well as the expectations of the subject matter experts in those fields, it is evident that this wish is simply not feasible. ML components are indeed software artifacts, but they are very special ones. As also stated by Lewis et al. 20, besides the data-dependent behavior of ML models, the necessity of detecting and responding to drift over time and timely capture of ground truth to inform retraining are among some of the unique characteristics of ML components that pose new challenges to software architects of ML-enabled systems and bring new quality attribute concerns (e.g., with respect to monitorability). Hence, we have to extend our software architecture frameworks if we want our software system architecture descriptions to include this vital aspect of ML-enabled systems, which is affecting the software systems’ behaviors and even structures in an unprecedented manner.
@@ 5.1 Identified stakeholders
Data scientists are responsible for analytics modeling. In fact, the task of developing methods for building efficient Data Analytics (DA) models (including ML models) to enable systems that can analyze data and learn from data lies at the core of data science. However, there is also a need for technologies regarding the deployment of DA models in products, services, and operational systems. This part is known as analytics operations. Data engineers are typically concerned with this part, which is also called Data Engineering (DE) 17. Together, DA (i.e., data science) and DE are called Data Engineering and Analytics (DEA) 24 or Data Science and Engineering (DSE) 25. The typical workflow comprises analytics modeling (e.g., training ML models) by data scientists, followed by the integration and deployment of the data analytics artifacts (e.g., the trained ML models) in the data analytics and ML components, as well as in the larger systems (i.e., ML-enabled systems) by data engineers in collaboration with software engineers, database engineers, system engineers, etc. Afterward, the system should be handed over to the operations team 16.
@@
5.2 Proposed viewpoints and views
We propose two new architecture viewpoint categories to frame the concerns of data scientists and data engineers in the architecture frameworks of systems, software, and enterprises. We call the new viewpoint categories analytics modeling (alternatively DA or data science) and analytics operations (or DE), respectively 17. Moreover, we propose adopting and adapting existing notations and model kinds to realize corresponding views for the new viewpoints, as well as new views for the viewpoints of other stakeholders communicating and collaborating with data scientists and data engineers.
According to our own knowledge about the state of practice in the field of DEA (or DSE), we envisioned the mathematical notations commonly used in DA and ML, as well as the graphical notation of PGMs 26, and the Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) 18 for ANNs, as potentially adequate candidates for providing the necessary model kinds that should serve the architecture views required for supporting the proposed new architecture viewpoints. In addition, we assumed that existing UML diagrams could be adopted for adapting the viewpoints of other stakeholders in order to enable their appropriate communication and collaboration with data scientists and data engineers. The empirical study through the survey questionnaire confirmed these assumptions and helped us devise new viewpoints and views using existing notations and model kinds, as illustrated in Table 5. The list of notations and model kinds in each row of the table is ordered based on the opinions of the survey participants concerning the suitability of each option for the specific purpose. The radar diagrams in Figures 2 and 3 show this for the collaboration of data scientists with their peers, whereas the ones in Figures 4 and 5 demonstrate this for the collaboration of data engineers with their peers. In fact, we asked the same question from each stakeholder group in two different ways: what is the best model kind and notation for their collaboration; and what is the most suitable model kind and notation for describing the system architecture from their viewpoint?
Table 5: Notations and model kinds supporting the viewpoints and views of ML-enabled architecture frameworks for framing and addressing stakeholder concerns with respect to ML
@@ 5.2 Proposed viewpoints and views
Data scientists with their peers i) mathematical notations, ii) charts, diagrams, or plots, iii) DFGs or CGs 18, iv) PGMs 26, v) data flow diagrams (or UML activity diagrams showing the flow of data rather than the flow of control), for example, for the data analytics pipeline
Data engineers with their peers i) data flow diagrams, ii) UML class diagrams, iii) DFGs or CGs 18, iv) Entity-Relationship (ER) diagrams, v) mathematical notations
@@
System designers, engineers, integrators, and network engineers with data scientists and engineers i) UML deployment diagrams, ii) data flow diagrams, iii) DFGs or CGs 18 augmented with physical (i.e., deployment) information, iv) UML class diagrams
Security experts with data scientists and engineers i) data flow diagrams, ii) UML deployment diagrams, iii) DFGs or CGs 18 augmented with physical (i.e., deployment) information, iv) ER diagrams, v) mathematical notations, vi) UML class diagrams
Safety and regulatory compliance engineers, data protection (privacy) officers, and ethics committees or boards with data scientists and engineers i) text documents, ii) data flow diagrams, iii) ER diagrams, iv) DFGs or CGs 18 augmented with physical (i.e., deployment) information, v) tables or matrices, vi) UML deployment diagrams
@@ 6 Threats to validity
7 Conclusion and future work
@@
One limitation of the present study was the absence of particular groups of stakeholders, such as the operations team (see 16), the maintenance managers, and quality assurance engineers. The two latter items were suggested by our online survey participants. In the future, these and other stakeholder groups can be studied. Also, the enhancement of architecture frameworks for other sub-disciplines of Artificial Intelligence (AI) beyond ML should be helpful.

### v3

@@ IV Research design and methodology
We conducted research in two steps: 1) We interviewed four selected domain experts to gain some insight into the problem domain and validate the interview questions. To this aim, we deployed the Qualitative Surveys (Interview Studies) empirical research method [21]. 2) We adopted the quantitative empirical research method, called Questionnaire Surveys [22] and carried out an online survey with 61 participants.
@@
We interviewed four experts from the authors’ networks from May to June 2021. The interviewee selection was based on convenience sampling (i.e., not random). The interviewee profiles were as described in Table II. The numbers in parentheses denote the frequencies. The interviews were semi-structured with open questions.
TABLE II: Summary of the demographic information of the interview partners
@@
We carried out a survey study from July to September 2021 to answer RQ1 and RQ2 above. The survey questionnaire was offered through a link (URL) as an online [^5] We used the open-source LimeSurvey software [23] on our own server., self-administered survey, with the option of fully anonymous participation in the study. However, we collected the IP addresses to prevent any possible redundant participation. The questionnaire had four sections and a total of 25 questions. The order of the questions in each section and the order of the choices in the case of multiple-choice questions would be set on a random basis for each participant. We had a total of 121 participants, out of which 60 participants did not answer the questionnaire at all. Therefore, we took the results of the remaining 61 participants into account. The selection process for the invitation of the subject matter experts to participate in this study was again based on convenience sampling (i.e., not random) through the authors’ networks, for example, by direct invitation of peers via email and sharing the URL on LinkedIn. Part of the participants’ demographic information is summarized in Table III (the numbers in parentheses denote the frequencies). The average participation time was 14 minutes, whereas the median was 11 minutes. All questions were optional to answer except for the two questions regarding the consent of the survey participants with respect to their anonymity and receiving a pre-print of the study in the future.
TABLE III: Summary of the survey participants’ demographic information
@@
TABLE IV: Summary of the demographic information of the interview partners in the second round of expert interviews
@@
V Enhancing architecture frameworks
Existing software system architecture frameworks fall short of supporting a holistic architecture description for ML-enabled systems. Some members of the Software Engineering (SE) community may argue that the ML-related functionality might be specified using existing functional, behavioral, or logical model kinds, and perhaps the structure of ML models might be specified using existing structural model kinds. However, looking at the reality and state of practice in ML, data science, and data engineering, as well as the expectations of the subject matter experts in those fields, it is evident that this wish is simply not feasible. ML components are indeed software artifacts, but they are very special ones. As also stated by Lewis et al. [20], besides the data-dependent behavior of ML models, the necessity of detecting and responding to drift over time and timely capture of ground truth to inform retraining are among some of the unique characteristics of ML components that pose new challenges to software architects of ML-enabled systems and bring new quality attribute concerns (e.g., with respect to monitorability). Hence, we have to extend our software architecture frameworks if we want our software system architecture descriptions to include this vital aspect of ML-enabled systems, which is affecting the software systems’ behaviors and even structures in an unprecedented manner.
@@ V-A Identified stakeholders
Data scientists are responsible for analytics modeling. In fact, the task of developing methods for building efficient Data Analytics (DA) models (including ML models) to enable systems that can analyze data and learn from data lies at the core of data science. However, there is also a need for technologies regarding the deployment of DA models in products, services, and operational systems. This part is known as analytics operations. Data engineers are typically concerned with this part, which is also called Data Engineering (DE) [17]. Together, DA (i.e., data science) and DE are called Data Engineering and Analytics (DEA) [24] or Data Science and Engineering (DSE) [25]. The typical workflow comprises analytics modeling (e.g., training ML models) by data scientists, followed by the integration and deployment of the data analytics artifacts (e.g., the trained ML models) in the data analytics and ML components, as well as in the larger systems (i.e., ML-enabled systems) by data engineers in collaboration with software engineers, database engineers, system engineers, etc. Afterward, the system should be handed over to the operations team [16].
@@
V-B Proposed viewpoints and views
We propose two new architecture viewpoint categories to frame the concerns of data scientists and data engineers in the architecture frameworks of systems, software, and enterprises. We call the new viewpoint categories analytics modeling (alternatively DA or data science) and analytics operations (or DE), respectively [17]. Moreover, we propose adopting and adapting existing notations and model kinds to realize corresponding views for the new viewpoints, as well as new views for the viewpoints of other stakeholders communicating and collaborating with data scientists and data engineers.
According to our own knowledge about the state of practice in the field of DEA (or DSE), we envisioned the mathematical notations commonly used in DA and ML, as well as the graphical notation of PGMs [26], and the Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) [18] for ANNs, as potentially adequate candidates for providing the necessary model kinds that should serve the architecture views required for supporting the proposed new architecture viewpoints. In addition, we assumed that existing UML diagrams could be adopted for adapting the viewpoints of other stakeholders in order to enable their appropriate communication and collaboration with data scientists and data engineers. The empirical study through the survey questionnaire confirmed these assumptions and helped us devise new viewpoints and views using existing notations and model kinds, as illustrated in Table V. The list of notations and model kinds in each row of the table is ordered based on the opinions of the survey participants concerning the suitability of each option for the specific purpose. The radar diagrams in Figures 2 and 3 show this for the collaboration of data scientists with their peers, whereas the ones in Figures 4 and 5 demonstrate this for the collaboration of data engineers with their peers. In fact, we asked the same question from each stakeholder group in two different ways: what is the best model kind and notation for their collaboration; and what is the most suitable model kind and notation for describing the system architecture from their viewpoint?
TABLE V: Notations and model kinds supporting the viewpoints and views of ML-enabled architecture frameworks for framing and addressing stakeholder concerns with respect to ML
@@ V-B Proposed viewpoints and views
Data scientists with their peers i) mathematical notations, ii) charts, diagrams, or plots, iii) DFGs or CGs [18], iv) PGMs [26], v) data flow diagrams (or UML activity diagrams showing the flow of data rather than the flow of control), for example, for the data analytics pipeline
Data engineers with their peers i) data flow diagrams, ii) UML class diagrams, iii) DFGs or CGs [18], iv) Entity-Relationship (ER) diagrams, v) mathematical notations
@@
System designers, engineers, integrators, and network engineers with data scientists and engineers i) UML deployment diagrams, ii) data flow diagrams, iii) DFGs or CGs [18] augmented with physical (i.e., deployment) information, iv) UML class diagrams
Security experts with data scientists and engineers i) data flow diagrams, ii) UML deployment diagrams, iii) DFGs or CGs [18] augmented with physical (i.e., deployment) information, iv) ER diagrams, v) mathematical notations, vi) UML class diagrams
Safety and regulatory compliance engineers, data protection (privacy) officers, and ethics committees or boards with data scientists and engineers i) text documents, ii) data flow diagrams, iii) ER diagrams, iv) DFGs or CGs [18] augmented with physical (i.e., deployment) information, v) tables or matrices, vi) UML deployment diagrams
@@ VI Threats to validity
VII Conclusion and future work
@@
One limitation of the present study was the absence of particular groups of stakeholders, such as the operations team (see [16]), the maintenance managers, and quality assurance engineers. The two latter items were suggested by our online survey participants. In the future, these and other stakeholder groups can be studied. Also, the enhancement of architecture frameworks for other sub-disciplines of Artificial Intelligence (AI) beyond ML should be helpful.